Subject: APO 3.5m -- DIS update and June users committee minutes

From: Suzanne Hawley

Submitted: Wed, 26 Jun 2013 09:47:52 -0700 (PDT)

Message number: 1314 (previous: 1313, next: 1315 up: Index)

As many of you know, and as described in the minutes below from the users 
committee meeting last week, DIS has been experiencing some issues so we 
took it offline for the past 10 days to investigate.  Thanks to excellent 
work by the APO staff, the instrument is now back in service and the 
problems appear to be largely resolved.  All DIS users for the next few 
weeks until shutdown are encouraged to examine their data closely and to 
send feedback to Suzanne, Russet and Bill (see cc line above for 
addresses).  We appreciate and depend on your help in monitoring data from 
all the instruments so that we can keep on top of any developing problems.

Suzanne Hawley
Director, ARC 3.5m Telescope
Apache Point Observatory


 		APO 3.5-m Users Committee Minutes, 6/17/13

Attending:  Suzanne Hawley, Mark Klaene, Scott Anderson, Michael Strauss, 
John Bally, Jon Holtzman, Mary Beth Kaiser, Remy Indebetouw, Bill 
Ketzeback, and Bruce Gillespie.

Absent:  Al Harper


User feedback and comments from institutional representatives:

   o Washington (Scott Anderson) - Scott had nothing new to report.

   o Chicago (Al Harper) - no report.

   o New Mexico State (Jon Holtzman) - Jon mentioned that in his reductions 
of DIS data from Jan/Feb, wavelength calibrations show wings in the 
spectral lines.  Suzanne said that this was previously noted, discussed, 
and steps are underway to remove the scattered light that is due to 
contamination in the dewar.  But one issue here is timeliness of getting 
feedback when there are instrumental problems, and Jon and Suzanne both 
stressed the continuing need for timely data inspection and communication 
between the remote observers and observatory staff.

   o Princeton (Michael Strauss) - Michael had nothing new to report.

   o Colorado (John Bally) - John had nothing new to report.

   o Virginia (Remy Indebetouw) - Remy had nothing new to report.

   o Johns Hopkins (Mary Beth Kaiser) - Mary Beth said that Hal Weaver was 
thankful for getting his recent comet observations during DD time, down to 
8 airmasses!


Discussion of telescope/instruments report:

Mark said that the hot and dry weather is abating, and we've seen clouds 
and a little moisture in the past weeks.  The main telescope issue has 
been occasional overheating of the drive amplifiers.  Agile is back in 
service, and is now working better but still not entirely robustly. 
SPIcam and TripleSpec are working fine, as is the echelle although the 
filter positioning software remains an issue.  The NIC-FPS ion pump was 
replaced.  As to the DIS red camera contamination, it was worked on last 
summer but it has since been determined that there may be a flaw in the 
design of its vacuum system -- we plan to warm the instrument this week 
and work on the problem.  Also under investigation is an increase in dark 
current seen in DIS data that correlates with the ambient temperature in 
the dome.

Suzanne said that the Users Committee members should inform their users 
about the problems with the DIS red camera (i.e., contamination, dark 
current), and consider switching to a blue-only science program until the 
problems are fixed if their DIS programs would be adversely affected by 
the red camera issues.  Jon asked if we could raise the temperature set 
point for the red camera to stabilize the dark current.  Mark said that 
the dark current at a higher set point would be unacceptable, and we can't 
drive the Cryotigers any colder.  Suzanne said that we will likely go into 
the dewar this summer; it would be helpful to talk with Jim Gunn and Jeff 
Morgan for some advice.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                5/2/13 through 6/12/13

0) Overview

Typical NM late spring weather, dry and hot, winds decreasing with a fair 
amount of dust which has affected observing to a small extent. Some 
localized cloud buildup and minor amounts of rainfall. Forest is in Stage 
II fire restrictions.

1) Telescope

The telescope performance has been nominal this period.  We are seeing 
amplifier overheats on the az and alt axes due to the hot temperatuires. 
These have been infrequent and caused little disruption to observing, 
however are troubling.  Upgraded amplifiers are available and we are 
researching them along with ways of improving airflow to the amplifiers.

2) Instruments

Low S/N was reported by a DIS user on the red side. We discovered that the 
dark current is much higher than normal. We also verified condensate 
inside the dewar.  Other users did not report issues and were still able 
to use DIS, but we have recently found that the dark current is highly 
correlated with the ambient temperature, so the effect on users varies 
from night to night.  We juggled the June schedule and will take DIS 
offline next week to investigate further.

Agile was inoperable this period. The camera was not taking images using 
the external trigger source. It was sent back to Princeton Instruments at 
the request of Anjum Mukadam. P.I. was unable to reproduce the failure but 
did update software and fixed a vacuum leak and cooling issue. The camera 
is back at APO and operating correctly most of the time. It appears to 
have an intermittent failure mode that we are still trying to 

SPIcam and TripleSpec operated nominally.

The echelle operated nominally, but the ecam filter wheel still fails when 
going from ND5 to ND1. This is a bug in the motor control software which 
has apparently been there since the instrument was delivered.  A permanent 
software fix is being investigated but for now avoid this movement.

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issue. NFS less than 9 is operational.  The ION pump failed 
completely and had to be replaced. No scheduled observing was affected.


Summer shutdown and Q3 schedule update:

Suzanne said that we had earlier planned for two separate shorter 
shutdowns this summer.  However, there have been delays in developing the 
software for migrating the TCC to Unix, so the second shutdown has been 
cancelled, and the first shutdown will be extended an additional week. 
Thus, the shutdown will run from 15 July - 8 August.  The Q3 observing 
schedule has been replanned accordingly.  There are a few extra bright 
open nights; these will be assigned soon.


AAS meeting feedback:

At the recent AAS meeting in Indianapolis, Suzanne said that there was 
much discussion about mid-sized telescopes.  In particular, NOAO is 
leaving the WIYN consortium, so the WIYN consortium held a session on 
possible new collaborations and the future landscape in general for 
4-meter-class telescopes, such as WIYN, ARC, Mayall, Steward, SOAR, 
Palomar, etc.  WIYN has separately asked if we would be interested in time 
sharing our telescope with them, and this is under active discussion. 
There followed an extended, and enthusiastic, discussion on time-sharing 
with various mid-sized telescopes.



[open from previous meeting]:

===> ACTION: Send new partner suggestions to Suzanne, Rene and Bruce.

 	Status:  Open, for as long as necessary.

===> ACTION:  Find users to assist with instrument documentation.

 	Status:  Open.  "Volunteers" are still reviewing documentation for 
feedback.  Additional help welcome.

===> ACTION:  Explore new instruments discussion with users.

 	Status:  Open.  No feedback yet.  Users Committee members should 
hold informal meetings with their institutional users.

[new from this meeting]:



Next meeting:  We will skip July since the telescope will be in shutdown. 
The next Users Committee phonecon will be on Monday 12 August at 8:30 AM 
Pacific Time.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1314 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to